OpenStudy (anonymous):

the second of two numbers is 3 less than twice the first their sum is 36. find the numbers

OpenStudy (anonymous):

13 and 23. If you have done algebra, you could use one variable as x and another as y. y=2x-3, x+y=36 so x+2x-3=36 or 3x=39, so x=13 and y =2(13)-3=26

OpenStudy (anonymous):

okay bmt this is a general principal you can use throughout math in your life. you need an equation for every unknown, in order to solve for both unknowns

OpenStudy (anonymous):

you have two unknowns as grodon909 mentioned x and y (the two unknowns)

OpenStudy (anonymous):

one equation you can see is addition of the two unknowns is equal to 36 so x + y = 36

OpenStudy (anonymous):

i know how to set them up i don't know how to do that problem that i said above

OpenStudy (anonymous):

so you NEED one more equation to solve for both x and y that equation is the one gordon909 said y = 2x - 3

OpenStudy (anonymous):

yeah its hard to understand how to set it up properly. you have to just take it step by step and write down each word into the equation

OpenStudy (anonymous):

Usually, what you want to do if you can is solve for one variable/unknown value (for example y was already in such a format) and use it in another equation (ie replace y=(some equation) for y in all cases that you see y).

OpenStudy (anonymous):

sometimes it seems slow and you wanna get it done. but its fast to get it right the first time then to have to redo it. i know that sounds corny but i found that is true everytime

OpenStudy (anonymous):

i give up